OpenStudy (anonymous):

is this series, \sum_0^infty(1/(2+sqrt(n))*e(int)) convergent?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Can you retype that please? :)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{1}{2+\sqrt{n}e^?}\]

OpenStudy (anonymous):

\[\sum_{0}^{\infty}{1/(2+\sqrt{n})e^{inx}}\]

OpenStudy (anonymous):

Ratio test?

OpenStudy (anonymous):

ratio =1

OpenStudy (anonymous):

it's a complex trigonometric series

OpenStudy (anonymous):

any thoughts?

OpenStudy (anonymous):

Hmmm...Not really. Lol. Other than having something like e^-inx=e^inx-2isin(nx) But I don't see how that would help at all.

OpenStudy (anonymous):

well thanks anyway :)